# Security and Quality in LLM-Generated Code: A Multi-Language, Multi-Model Analysis

### v1

@@ 3.6.1 Manual Method
The automated scripts were designed and written for each programming language to perform syntactic checks. For compiled languages such as C, C++, and Java, each source file was compiled, and the results of these compilation attempts were recorded. For interpreted languages such as Python, syntax validation was performed using Python’s native syntax-checking functionality.
The results were documented in a comprehensive matrix with 200 rows (representing the prompts) and 20 columns (corresponding to 5 LLMs and 4 languages). Each cell in the matrix was annotated with a binary value indicating syntactical validity and successful code compilation. This methodology facilitated the identification of potential issues, such as syntax errors or missing imports, enabling the resolution of missing library imports or the exclusion of problematic code before proceeding to semantic analysis.
Semantic and Functional Correctness. We evaluated the semantics of the generated code after verifying and reviewing all syntactic errors. This phase involved assessing whether the code produced by each LLM accurately implemented the logic and functionality specified in the task prompts.
Initially, our goal was to develop a single unit test file per language for each task prompt that could be universally applied to all LLM-generated solutions. However, during testing, we observed variations in the function signatures produced by different LLMs. Differences in parameter types, return values, function names, and parameter order made it impractical to use a single unit test file for all code generated by the selected LLMs for the same task.
To address this challenge, we created custom unit test files for each individual code file. Each test file contained ten unit test cases specifically designed to evaluate the correctness of that particular LLM output. For example, if an LLM-generated function had a unique signature—such as one function using an array data type as input and another using a vector data type–the unit test cases were adapted accordingly to match the expected input/output format.
Together, 4,000 different unit test files are required, 200 test files for each programming language per LLM. The results of these unit tests were documented using another matrix with the same matrix dimensions referenced in compilation-time error validation. Each unit test file receives a score ranging from 0% (all tests failed) to 100% (all tests passed), depending on how many test cases were successfully executed. This process is implemented by three developers with one round of review cycle for the written unit testing files. Through this process, we were able to measure the semantic accuracy of the generated code, providing a comprehensive insight into how effectively the LLMs comprehended the prompt specifications.
In summary, the evaluation faced key challenges, including inconsistencies in function signatures across LLMs, which hindered standardized unit testing. We addressed this by creating individualized unit test files, increasing complexity but ensuring accuracy. Additionally, variations in code quality and completeness required manual adjustments, such as adding missing imports, to enable proper compilation and testing.
3.6.2 Tool-Based Method
There are three evaluation metrics generated using the static code scanning tool.
Static Features. These features for codes generated by the LLMs are collected using SonarQube tool [27] which supports all selected languages in this study. These features are:
  1. Lines of Code (LoC) measures the program’s lines of code, excluding whitespace. LoC is a predictive metric used to evaluate effort and maintainability.
  2. Cyclomatic Complexity (CyC) calculates code complexity using a control flow graph (CFG). With $E$ as edges, $N$ as nodes, and $Q$ as connected components, CyC is computed as $M=E+2Q-N$ [35].
  3. Cyclomatic Complexity Density (CCD) measures how cyclomatic complexity spreads across the codebase. With $cl$ as the total code lines, CCD is calculated as $Md={(E+2Q-N)}/{cl}$ [36].
  4. Cognitive Complexity (CoC) measures the difficulty of understanding code [37]. CoC considers structure like control flow and nesting, using $C=C_{\text{base}}+\sum_{i=1}^{n}{nc}$, where $nc$ represents increases due to nesting and conditionals.
Software Quality Attributes. For our analysis, we assess the software quality of the generated code using SonarQube [27], a well-known tool for code quality inspection. Evaluate the source code on the basis of multiple quality metrics. Our evaluation focuses on four software quality indicators: ① reliability, ② security, ③ maintainability, ④ security hotspots. In total, we examined 97,412 lines of code generated in four languages by five different LLMs.
  1. Reliability measures how well the code operates under predefined conditions. The tool identifies bugs that may cause errors or unpredictable behavior. By finding and fixing these bugs, we ensure adherence to best practices and mitigate potential run-time issues. We analyzed bug density in the produced code to confirm high reliability, helping to prevent execution problems.
  2. Security ensures the code is free from vulnerabilities that malicious actors could exploit [38, 39]. The tool highlights security vulnerabilities, such as improper input handling or weak encryption, which may lead to breaches. We evaluated the number and severity of these vulnerabilities to ensure that the generated code meets modern security standards, protecting sensitive data and preventing unauthorized access.
  3. Maintainability refers to the ease of understanding, modifying, and extending the code over time. The tool detects code smells, indicating inefficient design choices that may hinder future development. We assessed the maintainability score by considering factors like code complexity and adherence to coding standards, ensuring the code remains flexible and manageable.
  4. Security Hotspots are areas of code that may not be vulnerabilities but are sensitive and could lead to security issues if mismanaged. These often involve security-critical functions such as authentication and data validation. The tool flags these areas for developer review, ensuring proper handling. We included an analysis of security hotspots to prevent the generated code from unintentionally introducing risks in critical sections.
Clean Code Attributes. When evaluating the attributes of clean code, we consider four primary dimensions: ① consistency, ② intentionality, ③ adaptability, ④ accountability. These characteristics offer a structure for evaluating the quality, readability, and maintainability of code produced from clean code practices and standards.
  1. Consistency assesses the formatting, naming conventions, and structural design of the code. It ensures that the generated code follows a standardized framework across different languages and prompts. This includes uniformity in spacing, indentation, and identifier casing. Maintaining consistency improves readability, facilitates collaboration, and reduces cognitive load during code review and maintenance.
  2. Intentionality evaluates the clarity and effectiveness of the code in fulfilling its purpose. It examines whether the generated code is logical, complete, and efficient, ensuring that it conveys its functionality without unnecessary complexity while maintaining coherent logic.
  3. Adaptability focuses on the ease in modifying the code. It measures whether the code is modular and structured to allow localized updates with minimal risk of introducing errors. Well-adaptable code maintains a clear separation of concerns, ensuring that each function or component serves a defined purpose, reducing overall complexity.
  4. Responsibility ensures adherence to ethical and professional standards. It evaluates whether the generated code complies with legal and licensing requirements, safeguards sensitive information, and uses inclusive language. This attribute helps maintain trust, professionalism, and ethical integrity in AI-generated code.
@@ 3.6.2 Tool-Based Method
Analysis. We analyze the evaluation results, concentrating on essential quality attributes such as validity, accuracy, security, reliability, maintainability and clean code attributes. A detailed analysis of the results is available in section 5.
@@ 4 Dataset
The dataset created to evaluate the proposed concept was meticulously crafted to effectively assess the code produced by the LLMs, such as those discussed above. 200 programming tasks were manually defined and classified to ensure comprehensive coverage of a wide range of programming concepts. The tasks were divided into different categories: ① problem-solving, general coding challenges that involve solving algorithmic problems, aimed at testing the logical and problem-solving capabilities of LLMs; ②algorithms, tasks that focus on implementing fundamental algorithms, such as sorting, searching, dynamic programming, and graph-based algorithms; ③ data structures, tasks designed to assess the correct usage, manipulation, and implementation of data structures such as arrays, linked lists, trees, graphs, and hashmaps; ④ secure coding, security-related prompts, carefully chosen based on the selected CWEs from MITRE Common Weakness Enumeration (CWE) beyond CWE Top 25, aimed at testing the LLM’s ability to generate code free from common vulnerabilities such as buffer overflows, and injection flaws; ⑤ concurrency and multithreading, tasks aimed at testing how well the generated code handles parallelism, thread synchronization, and race conditions; ⑥ programming and system, problems requiring the manipulation of file handling, database operations, networking, and error handling.
The prompts were chosen carefully and designed to ensure their solvability in various programming languages, some of these were derived from code challenge websites and dataset [40, 41, 42, 43]. This approach facilitated the assessment of LLM performance in Python, Java, C++, and C, while preserving uniformity in problem complexity and intent across all languages and LLMs.
@@
Features. The dataset consists of several attributes to guide the evaluation, including: ① task number, which refers to unique identifier assigned to each question; ② prompt title, which refers to brief title that summarizes the problem statement; ③ description, a detailed description of the task, outlining the problem to be solved; ④ hints, which refers to instructions that provided to AI model to guide the generation of the code; ⑤ solutions, represent the code generated by the LLM for each programming language along with the name of the model used; ⑥ source, which refers to the origin of the task, whether manually created or adapted; ⑦ test cases, which refers to 10 test cases per prompt were written to evaluate the semantic of the generated code; ⑧ tags, which refers to labels assigned to each task to facilitate the process of filtrating, categorization, and statistical analysis; ⑨ comments, which refers to notes made by reviewers to document the evaluation process, including issues encountered during the generation process.
Purpose and Scope. The dataset provides a comprehensive resource for evaluating LLM-generated code across multiple languages. It consists of 200 unique programming challenges, each solved by five LLMs in four languages, yielding 4,000 records. Each record represents a solution generated by a specific AI tool for a given task in a particular language, enabling detailed performance analysis.
Creation Guidelines. Creating the dataset followed a strict guideline to ensure that the generated data set refers to a high-quality dataset. The main guidelines are as follows:
✧ Task Design. Each programming task is created to provide a clear statement and make sure that the tasks created cover different topics, such as algorithms, data structure, and CWE-based questions.
✧ Task Distribution. The generated tasks were categorized into different topics to facilitate the evaluation of the AI-generated code in different problem domains.
✧ AI Tools Comparison. Different solutions were proposed by different LLMs (claude-3.5, gemini-1.5, codestral, GPT-4o, llama-3) in four programming languages (Python, Java, C++, and C).
✧ Partial Solutions. Each of the generated codes was evaluated; if LLMs fail to generate a complete solution for a specific problem or provide only partial code, this was documented. In this case, a partial score is given based on the number of passed test cases.

### v2

@@ 3.5.1 Semantic Soundness
For each programming language, automated scripts were created to perform syntax checks and were reviewed by one of the software engineers and the second author. For compiled languages (C, C++, and Java), each source file was compiled and the outcome recorded. For interpreted languages like Python, syntax validation was conducted using Python’s built-in checker. Results were documented in a matrix of 200 rows (tasks) by 20 columns (5 LLMs $\times$ 4 languages), where each cell contained a binary value indicating syntactic validity or successful compilation. This process enabled early identification of issues such as syntax errors or missing imports, allowing for the correction or exclusion of problematic samples prior to semantic analysis.
Semantic Correctness Using Unit Tests. We evaluated semantic correctness after resolving all syntactic errors by checking whether each generated program implemented the intended logic specified in the task prompt.
The initial goal was to create a single reusable unit test per task and language. Natural language prompts without fixed function signatures led to substantial variation in generated code, including differences in function names, parameters, return values, and I/O handling. This variability reflects realistic LLM use but prevented uniform test reuse.
We addressed this variability using a four-phase unit testing procedure. First, for each of the 200 tasks, we defined a language-agnostic specification describing intended behavior, input domains, and test scenarios, including normal, boundary, edge, and invalid cases. Reference implementations produced expected outputs. Second, after code generation, we analyzed each program’s interface, including function signatures, parameter types and order, output mechanisms, and data structures (e.g., Python lists versus std::vector in C++), which ruled out a unified test suite.
Third, we implemented per-sample unit tests tailored to each program’s structure while preserving the shared logical specification. All tests used the same inputs and expected outputs across languages and models, adapting only syntax and data types. Assertions targeted return values or captured standard output based on program behavior. Fourth, two professionals manually reviewed each test for correctness, and the first and second authors re-checked a representative subset across languages and task categories. This process enabled consistent and fair evaluation of semantic correctness despite structural differences in generated code.
In total, we created 4,000 unit test files, corresponding to 200 tests per programming language per LLM. Unit test outcomes were recorded in a matrix with the same dimensions used for compilation-time error analysis. Each unit test received a score from 0% (no tests passed) to 100% (all tests passed), based on the proportion of successful test cases. Two developers executed the testing, and the first and second authors conducted a single review cycle of the unit test files. This process quantified semantic correctness and reflected how accurately LLMs interpreted task specifications. The evaluation posed challenges, including inconsistent function signatures across LLM outputs, which prevented standardized testing. We addressed this by creating per-sample unit tests, increasing implementation effort but ensuring correctness. Additional variability in code quality and completeness required limited manual fixes, such as adding missing imports, to enable compilation and testing.
Beyond semantics, we analyzed code coverage on a representative subset of tasks using JaCoCo [37] for Java and comparable tools for other languages. Each task included up to 10 unit tests covering standard cases and edge conditions. Coverage analysis complemented correctness checks by assessing test thoroughness. Average line coverage across sampled tasks exceeded 75%, indicating that most logic paths in the generated code were exercised. This coverage level supports the robustness of test design and evaluation.
3.5.2 Complexity
Understanding code complexity is essential for assessing program maintainability and readability. Higher complexity increases the likelihood of bugs, complicates testing, and lowers developer productivity. In this work, we analyze established complexity metrics to capture various structural and cognitive aspects of difficulty in LLM-generated code. Lines of Code (LoC)measures the program’s lines of code, excluding whitespace. LoC is a predictive metric used to evaluate effort and maintainability. Cyclomatic Complexity (CyC)calculates code complexity using a control flow graph (CFG). With $E$ as edges, $N$ as nodes, and $Q$ as connected components, CyC is computed as M=E+2QNM=E+2Q-N [38]. Cyclomatic Complexity Density (CCD)measures how cyclomatic complexity spreads across the codebase. With $cl$ as the total code lines, CCD is calculated as Md=(E+2QN)/clMd={(E+2Q-N)}/{cl} [39]. Cognitive Complexity (CoC)measures the difficulty of understanding code [40]. CoC considers structure like control flow and nesting, using $C=C_{\text{base}}+\sum_{i=1}^{n}{nc}$, where $nc$ represents increases due to nesting and conditionals.
3.5.3 Quality and Security
We evaluate software quality and security using SonarQube [28] and CodeQL [36], two widely adopted static analyzers supporting all target languages. SonarQube is used to assess non-vulnerability quality dimensions, while CodeQL provides query-based validation of vulnerabilities. Consistent with prior work, we manually reviewed all reported vulnerabilities and validated a stratified 20% sample across languages and LLMs to mitigate false positives and contextual limitations. Our evaluation covers four indicators: reliability, security (vulnerabilities), maintainability, and security hotspots. In total, the analysis spans 97,412 lines of code across four languages by five LLMs. ① Reliability assesses correct operation under predefined conditions using reported bug findings. ② Security captures exploitable weaknesses, quantified by the number and severity of detected vulnerabilities. ③ Maintainability measures ease of understanding and modification based on code smells and complexity indicators. ④ Security Hotspots identify security-sensitive code regions requiring careful handling, such as authentication, authorization, and input validation.
3.5.4 Clean Code
We evaluate clean code in LLM-generated outputs using four dimensions derived from SonarQube’s static analysis: consistency, intentionality, adaptability, and responsibility.
  1. Consistency assesses formatting, naming conventions, and structural uniformity across languages and prompts, including spacing, indentation, and identifier casing. Consistent code improves readability and reduces review and maintenance overhead. SonarQube captures consistency issues via code smells such as improper naming, unused imports or dead code, and inconsistent indentation or braces, which are aggregated under maintainability due to their impact on code clarity.
  2. Intentionality evaluates whether the code clearly and efficiently implements the intended functionality without unnecessary complexity. SonarQube captures this attribute through cognitive complexity metrics and code smells that penalize deep nesting, large switch blocks, unclear logic, and bloated methods, directly contributing to maintainability.
  3. Adaptability measures ease of modification through modular design and separation of concerns, enabling localized changes with minimal risk. SonarQube assesses adaptability using function length, parameter count, and lack of modular decomposition (e.g., monolithic blocks), which affect maintainability and may trigger security hotspots when sensitive logic lacks encapsulation.
  4. Responsibility reflects compliance with professional and ethical standards, including secure data handling and avoidance of high-risk practices. We evaluate responsibility through detected vulnerabilities (e.g., hard-coded credentials), security hotspots (e.g., unchecked inputs or return values), and use of deprecated or insecure APIs, spanning both security and maintainability dimensions.
@@ 3.5.4 Clean Code
Analysis. We analyze the results across key quality attributes, including syntactic validity, semantic correctness, security, reliability, maintainability, and clean code properties. Detailed results appear in section 5.
To improve methodological transparency, Figure 1 summarizes the complete Quality Control (QC) pipeline applied to each task, covering dataset design, code generation, compilation, semantic testing, and static analysis.
1) Dataset. Dataset construction, tagging, difficulty labels, language-neutral behavioral specification and test scenarios.
QC. tasks, tags, and specs were peer-reviewed by two (professional) software engineers working alongside the first/second authors of this study. (Sec. 4.1, Sec. 3.5.1).
2) LLM code generation. For each task and language, five LLMs generate code under scripted prompts, outputs stored in structured repositories.
QC. scripted calls, fixed prompts, and traceable file naming ensure reproducibility, the quantity of generated programs was verified again by the first author to ensure that no LLM-produced output was missed (Sec. 3.4).
3) Compilation. Language-specific scripts compile/parse each file, compilation matrix recorded. Minor, consistent fixes (imports/includes) applied and recorded when aligned with the prompt.
QC. compilation failures and fixes inspected, logged, and reviewed by one of the software engineers and the second author. (Sec. 3.5.1).
4) Program unit tests. Using shared unit test specification, one unit test file per program, adapted to its interface (signatures, parameter order, data structures, return vs. print) while keeping identical logical cases.
QC. each test file was manually reviewed by two professionals to verify that it exercised the intended functional behavior, and a representative subset across languages and task categories was re-checked by the first and second authors (Sec. 3.5.1, App. E).
5) Execution and semantic scoring. Language-specific environment is used to execute all tests, per-program 0 - 100% semantic score computed from passed cases, coverage sampled for selected tasks.
QC. the process was carried out by two developers, with one review cycle for the written unit test files conducted by the first and second authors of this work. Summary matrices are inspected for outliers, unexpected failures trigger manual re-check of tests and specification (Sec. 3.5.1).
6) Static analysis. SonarQube evaluates reliability, security, maintainability, and hotspots, in addition to CodeQL using official query packs.
QC. unified configuration across languages, reports exported for manual inspection (Sec. 3.5.3).
7) Cross-tool comparison and reporting. Each SonarQube-reported vulnerability was manually verified for validity and cross-checked against CodeQL at the same file and location, after which per-language and per-model tables and matrices were compiled.
QC. discrepancies documented in tables, and supplementary spreadsheets in the replication package (Sec. 5, Sec. 7).
Figure 1: End-to-end quality-control, covering task design, code generation, compilation, semantic testing and scoring, static analysis, cross-tool comparison, and reporting.
@@ 4 Dataset
To advance the evaluation of the quality and security of LLM code generation, we developed a dataset that reflects practical and security-relevant challenges in real-world programming domains. Compared to previous work, the dataset offers several advantages: it is created manually, language-independent, and tailored to evaluate LLM-generated code across multiple dimensions, including code correctness, maintainability, reliability, and security. Our goal is to provide the research community with a robust foundation for testing, comparing, and improving LLMs in terms of functional correctness, code quality, and security-oriented scenarios by integrating tasks rooted in known vulnerabilities (e.g., CWEs) and actual programming practices, thereby enhancing their overall utility.
What distinguishes our dataset is both its breadth, spanning 200 tasks across diverse categories, and its depth, with each task annotated using expert-reviewed tags and accompanied by unit tests designed to evaluate LLM-generated code. The dataset is model-agnostic, supports four programming languages, and adheres to secure coding best practices. Thus, it functions as benchmarking resource and a flexible training and diagnostic tool for researchers and practitioners engaged in LLM-assisted code development.
4.1 Dataset Construction
Our dataset construction methodology consists of four major steps: functional categorization of tasks, task sources and cross-language adaptation, task validation and peer review, and task tagging and redundancy reduction. In the following, we review those steps in detail.
Functional Categorization. We first organized the tasks into functional categories, reflecting both foundational programming skills and real-world software security concerns:
  1. Secure Coding: tasks grounded in known software weaknesses based on MITRE’s CWE taxonomy, including issues such as injection flaws, cryptographic misuse, and insecure data handling;
  2. Data Structures and Algorithms: problems focusing on core computing concepts such as array manipulation, recursion, sorting, searching, and graph traversal;
  3. Parsing and Validation: tasks requiring careful input processing, constraint enforcement, and boundary checking to assess robustness and correctness;
  4. Networking: scenarios that emulate networked systems, including authentication errors and protocol misuse;
  5. Mathematics and Logic: challenges designed to evaluate numerical reasoning, precision, and logical flow;
  6. Programming Systems and Utilities: utility tasks involving file handling, scripting, system configuration, and general-purpose development workflows;
  7. Concurrency and Synchronization: tasks targeting race conditions, thread safety, and synchronization issues in parallel execution contexts.
Task Sources and Cross-Language Adaptation. We sourced tasks from coding challenge platforms, including LeetCode [41], Edabit [42], and Codewars [43], as well as from vulnerability databases [44] and academic datasets [22]. When suitable tasks were unavailable, we designed additional tasks based on realistic software engineering scenarios. We formulated each task in neutral language to support consistent implementation in Python, Java, C++, and C, enabling fair cross-language comparison. Tasks originally defined for a single language, such as Python tasks from SecurityEval, were rewritten for cross-language applicability without changing core logic, intent, or security focus.
Our design choice of concise, self-contained programming tasks, adopted partially from popular coding challenge platforms, is consistent with recent large-scale evaluations of LLM-generated code on tasks from such sources, where problems with clear specifications and automated tests are used to enable systematic comparison across models [45, 46, 47, 48]. In contrast to these works, our benchmark additionally incorporates security-focused tasks and supports four programming languages, while preserving the same style of well-specified, unit-testable problems. Table V summarizes the distribution of tasks by source and categorizes them by difficulty level. Only a small subset of the 200 is labeled Easy, while the large majority fall into the Medium or Hard categories. This distribution is intentional and reflects our focus on non-trivial problems that require reasoning about real-world cases, input validation, and security-relevant behavior rather than purely direct tasks.
Task Validation and Peer Review Process. To ensure the quality, clarity, and reliability of each task in the dataset, we implemented a structured peer review process involving two (professional) software engineers working alongside the second author of this study. The reviewers systematically assessed each task for clarity, difficulty level, and solvability. Their feedback was discussed and incorporated through multiple review cycles, with any ambiguities, inconsistencies, or edge cases resolved collaboratively.
Unit test is used to verify task correctness, where we established a unified unit test specification applied consistently across all studied programming languages. This specification defined identical inputs, expected outputs, and coverage scenarios, with only language-specific syntax that varied between implementations. An illustrative example of these test scenarios is provided in Appendix E. This unified method ensures that variations in results are attributable to differences in code generation quality rather than inconsistencies in validation, thus providing a fair and controlled basis for comparing LLM-generated codes across languages.
Table V: Task distribution by source and difficulty: Easy, Medium, and Hard.
  Source Description Difficulty Total
E M H
  Manual Custom tasks written by the first author of this research 3 117 3 123
Code challenge website Adopted from Codewars [43], LeetCode [41], Exercism [49], and Edabit [42] focused on algorithms, data structures and system design 14 43 7 64
Security dataset Adopted from SecurityEval [22] and CodeQL [50] focused on code security 0 13 0 13
Task Tagging and Redundancy Reduction. Each task was manually reviewed and tagged by at least two reviewers according to its functionality and any implicit security objectives. Tasks covering multiple concepts were assigned multiple tags. This approach enables the evaluation of LLM performance on both narrowly focused and multi-dimensional tasks. To avoid semantic redundancy, we conducted a manual similarity analysis across task prompts and implementations, removing tasks that exhibited significant overlap in logic or structure. In future work, we plan to incorporate automated semantic similarity techniques to further validate dataset diversity and reduce potential bias.
4.2 Dataset Features and Improvements
@@
The dataset consists of several attributes to guide the evaluation, including: ① task number, which refers to unique identifier assigned to each question; ② prompt title, which refers to brief title that summarizes the problem statement; ③ description, a detailed description of the task, outlining the problem to be solved; ④ hints, which refers to instructions that provided to AI model to guide the generation of the code; ⑤ solutions, represent the code generated by the LLM for each programming language along with the name of the model used; ⑥ source, which refers to the origin of the task, whether manually created or adapted; ⑦ test cases, which refers to 10 test cases per prompt were written to evaluate the semantic of the generated code; ⑧ tags; labels assigned to each task to facilitate filtrating, categorization, and statistical analysis; ⑨ comments, which refers to notes made by reviewers to document the evaluation process, including issues encountered during the generation process.
To improve interpretability and enable more focused evaluation, we consolidated the original set of task tags into a structured set of seven semantic categories. Tags were initially assigned through manual annotation based on detailed task analysis; however, as the dataset expanded, the tag set exhibited redundancy, inconsistent phrasing, and varying levels of specificity. To address this, we applied a semantic grouping strategy by clustering related tags under broader functional themes. For instance, technical concepts such as arrays, sorting, graphs, and recursion were unified under Data Structures and Algorithms, reflecting core computational topics. Likewise, all tags referencing Common Weakness Enumeration (CWE) identifiers, including CWE-20, CWE-89, and CWE-787, were categorized under Secure Coding (CWEs), encompassing tasks focused on recognizing and addressing software vulnerabilities. Table VI shows how tasks are distributed according to tag categories, achieved by combining similar tags into larger functional groups. A task may belong to one or several categories.
To illustrate dataset structure and content, Appendix B presents representative tasks from each programming category. These examples reflect the range of real-world scenarios in the benchmark and contextualize the evaluation by showing the scope and complexity of the prompts. We release the complete dataset as part of the replication package, including task descriptions, difficulty labels, unit tests, and model outputs, enabling detailed inspection and reproduction or extension of the evaluation.
Table VI: Distribution of tasks by tag categories after grouping related tags into broader functional areas. One task might fall into one or multiple categories.
Category Task Count
Secure Coding (CWEs - All CWE identifiers grouped) 102
Data Structures and Algorithms 86
Parsing and Validation 54
Networking 48
Mathematics and Logic 26
Programming Systems and Utilities 18
Concurrency and Synchronization 6
4.3 Secure Coding Coverage via CWE Grouping
To assess model behavior on security-critical tasks, we integrated 86 unique CWE identifiers into our dataset, as noted in the task tagging. Rather than presenting these individually, we grouped them into higher-level vulnerability domains based on their semantic alignment and the MITRE taxonomy. These categories include Input Validation (i.e.,, CWE-20, CWE-707), Injection (i.e.,, CWE-89, CWE-78), Buffer Errors (i.e.,, CWE-119, CWE-787), Access Control Issues, Cryptographic Issues, and others. This organization provides a clearer view of the dataset’s security coverage and enables targeted evaluation of model resilience against specific classes of vulnerabilities. Table VII presents the distribution of tasks across these security domains, and the full list of CWEs is available in our project repository.
Table VII: MITRE-aligned categorization of representative CWEs in the dataset.
CWE Category Representative CWEs
Input Validation CWE-20, CWE-707, CWE-642
Injection CWE-89, CWE-78, CWE-74
Buffer Errors CWE-119, CWE-787, CWE-125
Access Control Issues CWE-285, CWE-639, CWE-287
Cryptographic Issues CWE-321, CWE-326, CWE-327
Information Exposure CWE-209, CWE-532
Resource Management Errors CWE-400, CWE-770
Memory Errors CWE-416